Differentiate the following with regard to x

y=2x-8/5-3x^2


1
Expert's answer
2021-07-06T12:06:47-0400

"y={2x-8\over 5-3x^2}"


let "u=2x-8\implies {du\over dx}=2"


let "v=5-3x^2 \implies {dv\over dx}=-6x"


In this case, we use the quotient rule of differentiation as follows


"{dy\over dx}={({du\over dx}\times v)-(u\times{dv\over dx})\over v^2}"


"={2(5-3x^2 )-(2x-8)\times(-6x)\over (5-3x^2)^2}"


"={2(5-3x^2 )+6x(2x-8)\over (5-3x^2)^2}"


"={10-6x^2 +12x^2-48x\over (5-3x^2)^2}"


"={10+6x^2 -48x\over (5-3x^2)^2}"